The sum of a number and two times a smaller number is 62. Three times the bigger number exceeds the smaller number by 116.
Answer by TimothyLamb(4379)

x + 2y = 62
3x - y = 116
---
put the system of linear equations into standard form
---
x + 2y = 62
3x - y = 116
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 42
y = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php